# Comparing Bayesian and Frequentist Inference in Biological Models: A Comparative Analysis of Accuracy, Uncertainty, and Identifiability

### v2

@@ 6.2 GLM model for Lung Injury
The GLM provides a well-identified baseline example in which the data support stable parameter recovery under both frameworks (Section 5). Both approaches capture the rise and decline in the lung-injury time series with QDF providing a better fit (Figure 11).
Figure 11: Fitting visualization for the GLM model, with lung-injury data, using QDF and BFF approaches. Left QDF, and Right BFF.
@@
In calibration performance, QDF achieves lower point-error metrics (MAE 9.42 vs. 11.85; MSE 152.80 vs. 412.29; WIS 5.75 vs. 7.15) while both methods attain high PI coverage (Table 16). This pattern is consistent with an identifiable, low-dimensional setting where optimization-based estimation can yield very sharp point fits without compromising interval calibration.
The Bayesian fit converged well for this dataset ($\hat{R}=1$ for all GLM parameters; Table 17), supporting reliability of the reported posterior summaries.
@@
Table 16: Performance metrics for the GLM model, with lung-injury data, using QDF and BFF approaches.
@@
Table 17: Convergence diagnostics for the GLM model, with lung-injury data, using QDF and BFF approaches.
@@
The cross-method parameter comparison reinforces that $(r,p)$ are tightly constrained in both frameworks, with only modest differences in uncertainty widths (Figure 12).
Figure 12: Parameter comparison for the GLM model, with lung-injury data, using QDF and BFF approaches.
Across the summary metrics, the same ranking is apparent: QDF yields lower MAE/MSE/WIS while PI coverage remains comparable (Figure 13).
Figure 13: Performance metrics comparison for the GLM model, with lung-injury data, using QDF and BFF approaches.
The parameter bounds used for applying the BFF and QDF for this model is provided in Table 18 and 19.
Table 18: Parameter bounds for the GLM model, with lung-injury data, using QDF and BFF approaches.
Parameter Lower Bound (LB) Upper Bound (UB)
$r$ 0.0 NA
pp 0.0 NA
$k$ 0.0 NA
Table 19: Parameter bounds for the GLM model, with lung-injury data, using QDF and BFF approaches.
@@ 6.3 GLM model for mpox epidemic in the USA
For the 2022 U.S. mpox epidemic, both methods reproduce the overall rise–fall pattern, indicating that the GLM captures the main outbreak trajectory under either inference framework (Figure 14). Parameter estimates are broadly consistent across methods, with modest differences in uncertainty widths (Table 20). For example, $p$ is tightly concentrated around $\sim 0.8$–0.85 for both methods, while $r$ differs more in its central estimate (BFF 1.95 vs. QDF 2.25), reflecting method-specific tradeoffs in fitting the growth and saturation phases.
Figure 14: Fitting visualization for the GLM model, with U.S. mpox data, using QDF and BFF approaches. Left QDF, and Right BFF.
Calibration-period metrics favor QDF on point-error measures (MAE 9.45 vs. 11.739; MSE 150.47 vs. 399.43; WIS 5.8 vs. 7.175), while both methods attain high PI coverage (100% for QDF vs. 97.8% for BFF; Table 21). The Bayesian fit converged cleanly ($\hat{R}=1$ for all GLM parameters; Table 22), supporting stability of the reported BFF intervals.
Table 20: Parameter estimates for the GLM model, with U.S. mpox data, using QDF and BFF approaches.
@@
The cross-method parameter comparison (Figure 15) and metric summary (Figure 16) reinforce these trends: parameter uncertainty is generally tight under both methods, and QDF attains lower MAE/MSE/WIS while maintaining near-nominal PI coverage.
Table 21: Performance metrics for the GLM model, with U.S. mpox data, using QDF and BFF approaches.
@@
Table 22: Convergence diagnostics for the GLM model, with U.S. mpox data, using BFF approache.
@@
Figure 15: Parameter comparison for the GLM model, with U.S. mpox data, using QDF and BFF approaches. We use a log scale for the x-axis.
Figure 16: Performance metrics comparison for the GLM model, with U.S. mpox data, using QDF and BFF approaches.
The parameter bounds used for applying the BFF and QDF for this model is provided in Table 23 and 24.
@@
Parameter Lower Bound (LB) Upper Bound (UB)
$r$ 0.0 NA
pp 0.0 NA
$k$ 0.0 30000
@@ 6.4 SEIUR model of COVID-19 cases in Spain during the first wave
For Spain’s first COVID-19 wave, both methods track the rise and decline in reported incidence (Figure 17), but the inferred uncertainties and error metrics differ in ways consistent with the SEIUR model’s limited identifiability under incidence-only observations (Section 5). Compared with QDF, BFF yields lower errors (MAE 294.57 vs. 352.30; MSE 210,754 vs. 311,054; WIS 211.30 vs. 250.46) with similar PI coverage (94% vs. 92%; Table 26), suggesting an advantage for Bayesian regularization in this partially observed, higher-dimensional setting.
Figure 17: Fitting visualization for the SEIUR model, with Spain’s first COVID-19 wave data, using QDF and BFF approaches. Left QDF, and Right BFF.
The parameter summaries (Table 25) differ most for $\rho$ and $q_{1}$, which are strongly coupled to unobserved compartments and therefore weakly constrained by incidence alone. This helps explain why the two frameworks can yield different central estimates despite producing similar fitted trajectories.
Convergence diagnostics indicate some remaining sampling difficulty for the Bayesian SEIUR fit (Table 27; several parameters have $\hat{R}>1.1$), which is expected in a coupled, weakly identified setting. Accordingly, we interpret SEIUR posterior summaries with appropriate caution.
@@
Table 25: Parameter estimates for the SEIUR model, with Spain’s first COVID-19 wave data, using QDF and BFF approaches.
Table 26: Performance metrics for the SEIUR model, with Spain’s first COVID-19 wave data, using QDF and BFF approaches.
@@
Table 27: Convergence diagnostics for the SEIUR model, with Spain’s first COVID-19 wave data, using BFF approach.
@@
The metric summary (Figure 19) reinforces the table-based ranking: BFF achieves lower MAE/MSE/WIS while PI coverage remains close to nominal for both methods.
@@
The parameter bounds used for applying the BFF and QDF for this model is provided in Table 28 and 29.
Table 28: Parameter ranges for the SEIUR model, with Spain’s first COVID-19 wave data, using BFF approach.
Parameter Lower Bound (LB) Upper Bound (UB)
$\beta_{0}$ 0.0 NA
$\beta_{1}$ 0.0 NA
$q1$ 0.0 NA
$\rho$ 0.0 NA
$\kappa$ 0.0 NA
$\gamma_{1}$ 0.0 NA
NN 0.0 47000000
Table 29: Parameter ranges for the SEIUR model, with Spain’s first COVID-19 wave data, using QDF approach.
@@
NN 50000 1000000000.0
@@ 7 Discussion
This study compares Bayesian and Frequentist inference across ecological, clinical, and epidemiological models, with a focus on how data availability, model structure, and identifiability shape inference performance. Across all examples, we find that neither framework universally outperforms the other. Instead, performance depends critically on the degree of structural identifiability, the presence of latent states, and whether the primary goal is accurate point prediction or reliable uncertainty quantification. Bayesian methods, implemented via BayesianFitForecast (BFF), consistently provide stronger uncertainty estimates and diagnostic tools, while Frequentist methods, implemented via QuantDiffForecast (QDF), often achieve lower point prediction errors with substantially lower computational cost when identifiability is strong.
@@
For the Lotka–Volterra model, inference performance varied markedly across observation scenarios. When both prey and predator populations were observed, both methods produced stable and plausible parameter estimates. QDF yielded slightly lower point errors for predator dynamics, while BFF achieved better uncertainty calibration for prey. When only a single population was observed, QDF continued to provide sharper point predictions, whereas BFF maintained superior prediction interval coverage. Bayesian convergence diagnostics indicated stable posterior sampling even under partial observation, confirming the reliability of Bayesian inference in these settings.
@@
Differences in parameter estimates across observation schemes reflect changes in practical identifiability. With full observation of both species, interaction and turnover parameters are more strongly informed by the data, reducing confounding but also exposing differences between inference philosophies. In this regime, Bayesian and optimization-based approaches may favor different trade-offs when fitting coupled dynamics, leading to larger discrepancies in some parameters. Under partial observation, the Lotka–Volterra system becomes more weakly identifiable, as multiple parameter combinations can reproduce similar marginal dynamics for the observed series while uncertainty is absorbed by the unobserved state. In this case, both methods are effectively constrained toward similar parameter regimes through regularization, bounds, or prior structure, which can make several parameter estimates appear closer despite increased uncertainty.
@@
These empirical findings closely follow the predictions of the structural identifiability analysis. When both prey and predator are observed, all Lotka–Volterra parameters are structurally identifiable, whereas observing only one species leads to non-identifiability unless initial conditions are known. This limitation affects both Bayesian and Frequentist approaches equally, underscoring that observability places fundamental limits on inference quality regardless of statistical framework.
@@
For the generalized logistic model applied to lung injury and mpox outbreaks, QDF consistently achieved better point forecast accuracy while maintaining similar or better uncertainty calibration. In the mpox case, QDF produced lower MAE, MSE, and WIS, while achieving near-perfect or perfect prediction interval coverage. Bayesian inference showed excellent convergence for all parameters, confirming well-sampled posteriors. Similar patterns were observed for lung injury. These results are explained by the strong structural identifiability of the generalized logistic model: the single-equation structure and direct observation of incident cases ensure that all parameters are identifiable, even when initial conditions are unknown. In this setting, optimization-based inference provides accurate estimates efficiently, while Bayesian inference adds computational overhead without clear gains in predictive accuracy.
@@
A contrasting pattern emerges for the SEIUR model applied to COVID-19 data from Spain. Here, BFF outperformed QDF across all metrics, including point accuracy and uncertainty calibration. The SEIUR model includes latent compartments and exhibits limited structural identifiability when only reported cases are observed and initial conditions are unknown. Structural identifiability analysis shows that most parameters and state variables are unidentifiable in this setting, with only a small subset remaining identifiable. Under these conditions, Bayesian inference benefits from prior information, which regularizes the ill-posed inverse problem and stabilizes both parameter estimates and forecasts. In contrast, deterministic optimization must search a high-dimensional and weakly constrained parameter space, leading to less reliable results.
@@
Taken together, these results show that structural identifiability sets hard limits on what can be learned from data. No statistical method can recover parameters that are structurally unidentifiable, and identifiability analysis should therefore precede model fitting. When identifiability is strong and data are rich, Frequentist methods are well suited for minimizing point forecast error and computational cost. When identifiability is weak, data are sparse or noisy, or models include latent states, Bayesian methods provide more reliable uncertainty quantification and stable inference.
@@
Several methodological choices also influence these results. The use of different software environments reflects methodological requirements rather than implementation convenience: BFF relies on Stan-based Bayesian inference supported in R, while QDF depends on deterministic optimization and repeated numerical ODE solutions, for which MATLAB provides highly optimized solvers. Both approaches use identical model structures, data inputs, and evaluation metrics, ensuring that performance differences arise from inference methodology rather than computational platform.
@@
In the practical identifiability analysis, initial conditions were treated as known and fixed to align with the structural identifiability setting and to isolate differences arising from inference methodology. Allowing initial conditions to be estimated would introduce additional confounding, particularly in partially observed systems, and would be expected to increase uncertainty and variability in parameter estimates. Both QDF and BFF support joint estimation of initial conditions, and examining the impact of this added uncertainty remains an important direction for future work.
@@
This study intentionally focuses on low-dimensional models to allow transparent interpretation and the application of differential-algebraic structural identifiability methods. While this enables a clean methodological comparison, it also imposes limitations. We assumed normal error models, analyzed systems independently, and considered idealized identifiability conditions. For larger and more realistic models, practical identifiability tools based on Fisher information or sensitivity analysis may be necessary, especially under noisy or incomplete observations. Similarly, while Hamiltonian Monte Carlo performs well for moderate parameter dimensions, higher-dimensional models may require alternative sampling strategies such as adaptive Metropolis or DRAM.
@@
Supplementary Material

### v3

@@ 6.2 GLM model for lung injury
The GLM provides a well-identified baseline example in which the data support stable parameter recovery under both frameworks (Section 5). Both approaches capture the rise and decline in the lung injury time series, with QDF providing a better fit (Figure 11).
Figure 11: Fitting visualization for the GLM model, with lung injury data, using the QDF and BFF approaches. Left: QDF, Right: BFF.
@@
In calibration performance, QDF achieves lower point-error metrics (MAE 9.42 vs. 11.85; MSE 152.80 vs. 412.29; WIS 5.75 vs. 7.15) while both methods achieve well calibrated PI coverage near the nominal 95% level (97.83% for BFF; 100% for QDF), noting that 100% reflects slightly overconservative intervals (Table 16). This pattern is consistent with an identifiable, low-dimensional setting where optimization-based estimation can yield very sharp point fits without compromising interval calibration.
The Bayesian fit converged well for this dataset ($\hat{R}=1$ for all GLM parameters; Table 17), supporting the reliability of the reported posterior summaries.
@@
Table 16: Parameter estimates for the GLM model, with lung injury data, using the QDF and BFF approaches.
@@
Table 17: Convergence diagnostics for the GLM model with lung injury data, using the QDF and BFF approaches.
@@
The cross-method parameter comparison reinforces that $(r,p)$ are tightly constrained in both frameworks, with only modest differences in their uncertainty widths (Figure 12).
Figure 12: Parameter comparison for the GLM model, with lung injury data, using the QDF and BFF approaches.
Across the summary metrics, the same ranking is apparent: QDF yields lower MAE, MSE, WIS while the PI coverage remains comparable (Figure 13).
Figure 13: Performance metrics comparison for the GLM model, with lung injury data, using the QDF and BFF approaches.
The parameter bounds used for applying the BFF and QDF for this model are provided in Tables 18 and 19.
Table 18: Parameter bounds for the GLM model with lung injury data, using the BFF approach.
Parameter Prior distribution Lower bound (LB) Upper bound (UB)
$r$ Normal $(0.5,0.5)$ T[0,][0,\ ] 0.0 NA
$p$ Normal $(0.5,0.5)$ T[0,][0,\ ] 0.0 NA
$k$ Normal $(2632,10)$ T[0,][0,\ ] 0.0 NA
Table 19: Parameter bounds for the GLM model with lung injury data, using the QDF approach.
@@
The histogram summaries (Figures 2627) present the posterior, or sampling, distributions of the GLM parameters $(r,k,p)$ under BFF and QDF. Each panel reports the mean, median, and 95% confidence interval (CI), enabling a direct comparison of parameter uncertainty between the two estimation frameworks. Overall, both methods yield relatively concentrated distributions, suggesting that the GLM parameters are well identified; for example, $p\approx 1.0\pm 0.02$.
@@ 6.3 GLM model for mpox epidemic in the USA
For the 2022 U.S. mpox epidemic, both methods reproduce the overall rise–fall pattern, indicating that the GLM captures the main outbreak trajectory under either inference framework (Figure 14). Parameter estimates are broadly consistent across methods, with modest differences in the uncertainty widths (Table 20). For example, $p$ is tightly concentrated around 0.80.8–0.85 for both methods, while $r$ differs more in its central estimate (BFF 1.95 vs. QDF 2.25), reflecting method-specific trade-offs in fitting the growth and saturation phases.
Figure 14: Fitting visualization for the GLM model, with US mpox data, using the QDF and BFF approaches. Left: QDF, Right: BFF.
Calibration-period metrics favor QDF on point-error measures (MAE: 9.45 vs. 11.739; MSE: 150.47 vs. 399.43; WIS: 5.8 vs. 7.175), while both methods achieves coverage near the nominal 95% level (97.8% for BFF; 100% for QDF), with QDF’s 100% indicating slightly overconservative intervals; Table 21). The Bayesian fit converged cleanly ($\hat{R}=1$ for all GLM parameters; Table 22), supporting the stability of the reported BFF intervals.
Table 20: Parameter estimates for the GLM model with US mpox data, using the QDF and BFF approaches.
@@
The cross-method parameter comparison (Figure 15) and metric summary (Figure 16) reinforce these trends: parameter uncertainty is generally tight under both methods, and QDF achieves lower MAE, MSE, WIS while maintaining near-nominal PI coverage.
Table 21: Performance metrics for the GLM model with US mpox data, using the QDF and BFF approaches.
@@
Table 22: Convergence diagnostics for the GLM model with US mpox data, using the BFF approache.
@@
Figure 15: Parameter comparison for the GLM model, with U.S. mpox data, using QDF and BFF approaches. We use a log scale for the x-axis in the $K$ panel.
Figure 16: Performance metrics comparison for the GLM model with US mpox data, using the QDF and BFF approaches.
The parameter bounds used for applying the BFF and QDF for this model are provided in Tables 23 and 24.
@@
  Parameter   Prior distribution   Lower bound (LB)   Upper bound (UB)
  $r$   Normal $(0.5,0.5)$ T[0,][0,\ ]   0.0   NA
  $p$   Normal $(0.5,0.5)$ T[0,][0,\ ]   0.0   NA
  $k$   Normal $(29751,10)$ T[0,][0,\ ]   0.0   30000
@@
The histogram summaries (Figures 2829) show the posterior distributions under BFF and the sampling distributions under QDF for the GLM parameters $(r,k,p)$ fitted to the mpox incidence data. For each parameter, the panels report the mean, median, and 95% confidence interval (CI), allowing direct side-by-side comparison of uncertainty between the two estimation frameworks. Overall, both methods produce tight, largely unimodal distributions, suggesting that the GLM parameters are well identified. In particular, the parameters such as $p$, $k$, and $r$ exhibit concentrated uncertainty summaries across the two approaches, though the exact figure labeling should be checked carefully before submission.
@@ 6.4 SEIUR model of COVID-19 cases in Spain during the first wave
For Spain’s first COVID-19 wave, both methods track the rise and decline in reported incidence (Figure 17), but the inferred uncertainties and error metrics differ in ways that are consistent with the SEIUR model’s limited identifiability under incidence-only observations (Section 5). Compared with QDF, BFF yields lower errors (MAE 294.57 vs. 352.30; MSE 210,754 vs. 311,054; WIS 211.30 vs. 250.46) with similar PI coverage (94% vs. 92%; Table 26), suggesting an advantage for Bayesian regularization in this partially observed, higher-dimensional setting.
Figure 17: Fitting visualization for the SEIUR model with Spain’s first COVID-19 wave data, using the QDF and BFF approaches. Left: QDF, and Right: BFF.
The parameter summaries (Table 25) differ most for $\rho$ and $q_{1}$, which are strongly coupled to unobserved compartments and are therefore weakly constrained by incidence alone. This helps explain why the two frameworks can yield different central estimates despite producing similar fitted trajectories.
The convergence diagnostics indicate some remaining sampling difficulty for the Bayesian SEIUR fit (Table 27; several parameters have $\hat{R}>1.1$), which is expected in a coupled, weakly identified setting. Accordingly, we interpret SEIUR posterior summaries with appropriate caution.
Table 25: Parameter estimates for the SEIUR model with Spain’s first COVID-19 wave data, using the QDF and BFF approaches.
@@
Table 26: Performance metrics for the SEIUR model, with Spain’s first COVID-19 wave data, using the QDF and BFF approaches.
@@
Table 27: Convergence diagnostics for the SEIUR model with Spain’s first COVID-19 wave data, using the BFF approach.
@@
The metric summary (Figure 19) reinforces the table-based ranking: BFF achieves lower MAE, MSE, WIS while PI coverage remains close to nominal for both methods.
@@
The parameter bounds used for applying the BFF and QDF for this model are provided in Tables 28 and 29.
Table 28: Parameter ranges for the SEIUR model with Spain’s first COVID-19 wave data, using the BFF approach.
  Parameter   Prior distribution   Lower bound (LB)   Upper bound (UB)
  $\beta_{0}$   Normal $(0.5,0.5)$ T[0,][0,\ ]   0.0   NA
  $\beta_{1}$   Normal $(0.5,0.5)$ T[0,][0,\ ]   0.0   NA
  $q1$   Normal $(0.5,0.5)$ T[0,][0,\ ]   0.0   NA
  $\rho$   Normal $(0.5,0.5)$ T[0,][0,\ ]   0.0   NA
  $\kappa$   Normal $(0.5,0.5)$ T[0,][0,\ ]   0.0   NA
  $\gamma_{1}$   Normal $(0.5,0.5)$ T[0,][0,\ ]   0.0   NA
Table 29: Parameter ranges for the SEIUR model with Spain’s first COVID-19 wave data, using the QDF approach.
@@
The histogram summaries (Figures 3031) show the posterior distributions under BFF and the sampling distributions under QDF for the SEIR model parameters. For each parameter, the panels report the mean, median, and 95% confidence interval (CI), allowing side-by-side comparison of uncertainty between the two estimation frameworks. The BFF histograms are comparatively broad for several rate parameters, reflecting partial identifiability from the observed data, whereas the QDF histograms tend to be narrower and more clearly unimodal, indicating greater stability in the corresponding estimates. Together, these results emphasize the role of parameter coupling and limited data information in shaping uncertainty, and they suggest that additional data or reparameterization could improve identifiability.
@@ 7 Discussion
This study compares Bayesian and frequentist inference across ecological, clinical, and epidemiological models, with a focus on how data availability, model structure, and identifiability shape the inference performance. Across all examples, we find that neither framework universally outperforms the other. Instead, performance depends critically on the degree of structural identifiability, the presence of latent states, and whether the primary goal is accurate point prediction or reliable uncertainty quantification. Bayesian methods, implemented via BFF, consistently provide stronger uncertainty estimates and diagnostic tools, while frequentist methods, implemented via QDF, often achieve lower point prediction errors with substantially lower computational cost when identifiability is strong.
@@
For the LV model, inference performance varied markedly across observation scenarios. When both prey and predator populations were observed, both methods produced stable and plausible parameter estimates. QDF yielded slightly lower point errors for predator dynamics, while BFF achieved better uncertainty calibration for prey. When only a single population was observed, QDF continued to provide sharper point predictions, whereas BFF maintained superior PI coverage. Bayesian convergence diagnostics indicated stable posterior sampling even under partial observation, confirming the reliability of Bayesian inference in these settings.
@@
Differences in parameter estimates across observation schemes reflect changes in practical identifiability. With full observation of both species, the interaction and turnover parameters are more strongly informed by the data, reducing confounding but also exposing differences between inference philosophies. In this regime, Bayesian and optimization-based approaches may favor different trade-offs when fitting coupled dynamics, leading to larger discrepancies in some parameters. Under partial observation, the LV system becomes more weakly identifiable, as multiple parameter combinations can reproduce similar marginal dynamics for the observed series while uncertainty is absorbed by the unobserved state. In this case, both methods are effectively constrained toward similar parameter regimes through regularization, bounds, or the prior structure, which can make several parameter estimates appear closer despite increased uncertainty.
@@
These empirical findings closely follow the predictions of the structural identifiability analysis. When both the prey and predator are observed, all LV parameters are structurally identifiable, whereas observing only one species leads to nonidentifiability unless the initial conditions are known. This limitation affects both the Bayesian and frequentist approaches equally, underscoring that observability places fundamental limits on inference quality regardless of the statistical framework.
@@
For the GLM applied to lung injury and mpox outbreaks, QDF consistently achieved better point forecast accuracy while maintaining similar or better uncertainty calibration. In the mpox case, QDF produced lower MAE, MSE, and WIS, while achieving PI coverage at or slightly above the nominal 95% level, indicating conservative uncertainty estimates rather than improved predictive accuracy. Bayesian inference showed excellent convergence for all parameters, confirming well-sampled posteriors. Similar patterns were observed for lung injury. These results are explained by the strong structural identifiability of the GLM: The single-equation structure and direct observation of incident cases ensure that all parameters are identifiable, even when the initial conditions are unknown. In this setting, optimization-based inference provides accurate estimates efficiently, while Bayesian inference adds computational overhead without clear gains in predictive accuracy.
@@
A contrasting pattern emerges for the SEIUR model applied to COVID-19 data from Spain. Here, BFF outperformed QDF across all metrics, including point accuracy and uncertainty calibration. The SEIUR model includes latent compartments and exhibits limited structural identifiability when only reported cases are observed and the initial conditions are unknown. Structural identifiability analysis shows that most parameters and state variables are unidentifiable in this setting, with only a small subset remaining identifiable. Under these conditions, Bayesian inference benefits from prior information, which regularizes the ill-posed inverse problem and stabilizes both parameter estimates and forecasts. In contrast, deterministic optimization must search a high-dimensional and weakly constrained parameter space, leading to less reliable results.
@@
Taken together, these results show that structural identifiability sets hard limits on what can be learned from data. No statistical method can recover parameters that are structurally unidentifiable, and identifiability analysis should therefore precede model fitting. When the identifiability is strong and the data are rich, frequentist methods are well suited for minimizing point forecast error and computational cost. When the identifiability is weak, the data are sparse or noisy, or the models include latent states, Bayesian methods provide more reliable uncertainty quantification and stable inference.
@@
Several methodological choices also influence these results. The use of different software environments reflects methodological requirements rather than implementation convenience: BFF relies on Stan-based Bayesian inference supported in R, while QDF depends on deterministic optimization and repeated numerical ODE solutions, for which MATLAB provides highly optimized solvers. Both approaches use identical model structures, data inputs, and evaluation metrics, ensuring that performance differences arise from the inference methodology rather than the computational platform.
@@
In the practical identifiability analysis, the initial conditions were treated as known and fixed to align with the structural identifiability setting and to isolate differences arising from the inference methodology. Allowing the initial conditions to be estimated would introduce additional confounding, particularly in partially observed systems, and would be expected to increase the uncertainty and variability in parameter estimates. Both QDF and BFF support joint estimation of the initial conditions, and examining the impact of this added uncertainty remains an important direction for future work.
@@
This study intentionally focuses on low-dimensional models to allow transparent interpretation and the application of differential algebraic structural identifiability methods. While this enables a clean methodological comparison, it also imposes limitations. We assumed normal error models, analyzed the systems independently, and considered idealized identifiability conditions. For larger and more realistic models, practical identifiability tools based on Fisher information or sensitivity analysis may be necessary, especially under noisy or incomplete observations. Similarly, while Hamiltonian Monte Carlo performs well for moderate parameter dimensions, higher-dimensional models may require alternative sampling strategies such as adaptive Metropolis or Delayed Rejection Adaptive Metropolis (DRAM).
Practical implications
Our results provide practical guidance for selecting between Bayesian and frequentist inference frameworks:
  • Use frequentist methods (QDF) when: (1) the data are abundant and of high quality, (2) the models are structurally identifiable, (3) computational speed is critical, (4) the primary objective is minimizing point forecast error, and (5) the model’s complexity is low to moderate. The GLM applications exemplify these conditions.
  • Use Bayesian methods (BFF) when: (1) the data are sparse, noisy, or partially observed, (2) the models contain latent states or unobserved compartments, (3) the structural identifiability is limited; (4) comprehensive uncertainty quantification is essential; (5) prior information is available to constrain inference; and (6) diagnostic assurance via convergence metrics is important. The SEIUR application exemplifies these conditions.
  • Improve both methods by: (1) increasing observational coverage (e.g., measuring both the prey and the predator rather than one species), (2) obtaining accurate estimates of the initial conditions, (3) incorporating auxiliary data sources to break the parameters’ nonuniqueness, and (4) using an SI-guided experimental design to ensure parameters of interest are identifiable.
When practical identifiability is limited or interval calibration is critical, BFF is advantageous because posterior sampling naturally captures the parameter and predictive uncertainty while providing convergence assurance (e.g., GLM $\hat{R}=1$; LV $\hat{R}\approx 1.01$). When the priority is minimizing point forecast error and the data are informative, QDF can be preferable (e.g., LV for prey only; GLM for mpox), though PIs may be under calibrated relative to BFF (e.g., the LV for prey only coverage was 90.48% vs. BFF’s higher coverage, which reflects more conservative uncertainty rather than improved predictive accuracy). Across all systems, adding informative observations improves both methods (LV for predator + prey vs. single-series), and SI analysis clarifies when further accuracy gains are impossible without additional data or constraints.
@@
Use of AI tools declaration
The authors declare they have not used artificial intelligence (AI) tools in the creation of this article.
Conflict of interest
The authors declare there is no conflict of interest.
Appendix materials